OpenStudy (anonymous):

Change the subject of the following formulas to the indicated subject. P = 2w + 2l w= C = π d d=

OpenStudy (anonymous):

\[2(w+l)=P\] \[w+l=P/2\] \[w=P/2-l\]

OpenStudy (anonymous):

w=(P/2)-l \[d=C/\pi\]